OpenStudy (anonymous):

solve each 1st degree equations with radicals (square root of x-8)=2

OpenStudy (anonymous):

\[\sqrt{x-8} =2\]

OpenStudy (anonymous):

simple square both sides

OpenStudy (saifoo.khan):

\[x-2 = 2^2\]

OpenStudy (anonymous):

Square both sides to get rid of the radical \[(\sqrt{x-8})^2 = 2^2\] Squaring a radical cancels out the radical leaving you with \[x - 8 = 4\] Solve for x \[x = 12\]

OpenStudy (anonymous):